how do you solve: the sum of 4 times lisa's age and 7 times janes age is 169. jane is 1 year more than twice as old as lisa is. find each of their ages
Answer by rfer(16322)

4x+7(2x+1)=169
4x+14x+7=169
18x=162
x=9
-------------
4*9+7(2*9+1)=169
36+126+7=169
-------------
L=x=9
A=2x+1=19